### v1

# Pranav Suryadevara Morris Plains, NJ, USA

Abstract
The growth of machine learning (ML) workloads has underscored the importance of efficient memory hierarchies to address bandwidth, latency, and scalability challenges. HERMES focuses on optimizing memory subsystems for RISC-V architectures to meet the computational needs of ML models such as CNNs, RNNs, and Transformers. This project explores state-of-the-art techniques such as advanced prefetching, tensor-aware caching, and hybrid memory models. The cornerstone of HERMES is the integration of shared L3 caches with fine-grained coherence protocols and specialized pathways to deep learning accelerators like Gemmini. Simulation tools like Gem5 and DRAMSim2 are used to evaluate baseline performance and scalability under representative ML workloads. The findings of this study highlight the design choices and anticipated challenges, paving the way for low-latency scalable memory operations for ML applications.
Index Terms:
RISC-V, Memory Hierarchy, Machine Learning, Shared L3 Cache, Tensor-Aware Caching, Advanced Prefetching, Gemmini, Gem5, DRAMSim2, ML Workloads, Memory Bottlenecks.
I Introduction
Machine learning (ML) models have become integral to modern computing, with applications spanning computer vision, natural language processing, and autonomous systems. These models, such as Convolutional Neural Networks (CNNs), Recurrent Neural Networks (RNNs), and Transformers, often require extensive computational resources and memory bandwidth. Consequently, optimizing memory hierarchies for ML workloads is critical to improving performance and efficiency [1, 2].
RISC-V, an open source Instruction Set Architecture (ISA), is gaining traction due to its flexibility and extensibility [3, 4]. However, the performance of RISC-V processors for ML workloads is hindered by memory bottlenecks, including latency, limited bandwidth, and lack of scalability [6]. The HERMES project aims to design a high-performance memory subsystem tailored to ML workloads on RISC-V architectures. Using techniques such as advanced prefetching [8], tensor-aware caching [9], and hybrid memory models [7], HERMES seeks to overcome these challenges and enhance the efficiency of ML computations.
This report presents a foundational study of HERMES, outlining the design considerations, initial simulation results, and challenges in achieving a scalable, low-latency memory hierarchy for ML workloads and evaluate the performance using simulation tools such Gem5 [10] and DRAWSim2[11].
II Literature Survey
II-A Memory Bottlenecks in ML Workloads
ML workloads, especially deep learning models, are characterized by large datasets and frequent memory access patterns. Studies have identified key challenges that impact the efficiency of memory subsystems:
High Latency
Tensor operations, such as matrix multiplications, require frequent access to large datasets stored in memory. The latency involved in fetching data from DRAM can significantly degrade performance if not mitigated by efficient caching and prefetching mechanisms [1, 2, 5].
Bandwidth Constraints
ML models, particularly those involving convolutions and recurrent operations, demand high memory bandwidth. The concurrent fetching of weights, activations, and gradients often saturates the available bandwidth, leading to bottlenecks in data transfer. This issue is exacerbated in systems with limited on-chip memory [6, 7, 12].
Scalability
As ML models grow in size and complexity, memory subsystems must scale to accommodate increasing data loads. Current architectures struggle to maintain performance as the dataset size exceeds cache capacity, requiring frequent access to slower off-chip memory [10, 15].
Addressing these bottlenecks requires innovative memory hierarchies that can balance latency, bandwidth, and scalability for different ML workloads.
II-B Existing Memory Solutions for ML
Various techniques have been developed to optimize memory systems for ML workloads. These solutions address specific aspects of memory bottlenecks and have been implemented in modern hardware architectures.
Tensor-Aware Caching
Tensor-aware caching improves data reuse by optimizing the cache layout for tensor operations. For example, Google’s Tensor Processing Unit (TPU) uses specialized caching mechanisms that align with the structure of tensor computations, reducing latency and increasing throughput [1]. This approach ensures that frequently accessed tensors remain in fast on-chip memory, minimizing DRAM accesses.
Prefetching Strategies
Advanced prefetching techniques predict data access patterns and load data into the cache before it is needed. Techniques such as Stride Prefetching and Machine Learning-Based Prefetching have proven effective in reducing latency for sequential and non-sequential data access patterns [8, 5]. For instance, Intel’s microprocessors incorporate sophisticated hardware prefetchers that dynamically adapt to workload behavior [9, 12].
Hybrid Memory Models
Combining DRAM with High-Bandwidth Memory (HBM) offers a balance between capacity and speed. HBM provides significantly higher bandwidth compared to traditional DRAM, making it suitable for ML workloads with high data transfer requirements [7, 16]. For example, NVIDIA’s GPUs use HBM to deliver the necessary bandwidth for training large neural networks [13].
These techniques have shown promise in improving memory performance for ML workloads, but their integration in RISC-V architectures remains limited.
II-C RISC-V-Based Accelerators
The open source nature of RISC-V has led to the development of several accelerators designed for ML workloads. These accelerators benefit from custom memory pathways and specialized hardware support for tensor computations.
Gemmini
Gemmini is a configurable deep learning accelerator for RISC-V systems. It provides efficient support for matrix multiplication and convolution operations, which are fundamental to ML workloads [6, 14]. The Gemmini design allows integration with custom memory pathways, facilitating efficient data transfer between caches and the accelerator.
Simulation Tools
Tools like Gem5 and DRAMSim2 are widely used to evaluate memory hierarchies and system performance. Gem5 offers a flexible platform for modeling CPU and memory interactions [10, 18], while DRAMSim2 provides detailed simulations of DRAM behavior [11, 17]. These tools are essential for testing and validating new memory subsystem designs in RISC-V architectures.
II-D Research Gaps
While existing solutions address specific aspects of memory optimization, there is limited research on integrating these techniques into a cohesive memory subsystem for RISC-V architectures specifically targeting ML workloads. Most current implementations focus on proprietary architectures, such as TPUs and GPUs, leaving a gap in the open-source RISC-V ecosystem [15, 13].
HERMES: A Unified Memory Solution
HERMES aims to fill this gap by:
  1. 1.
    Combining shared L3 caches, hybrid memory models, tensor-aware caching, and advanced prefetching into a unified memory hierarchy.
  2. 2.
    Optimizing memory pathways for Gemmini and similar RISC-V accelerators [6].
  3. 3.
    Providing a scalable and efficient solution for deep learning models on RISC-V platforms.
III Problem Statement
Current RISC-V memory hierarchies face significant challenges in meeting the performance requirements of ML workloads due to several interrelated issues:
  1. 1.
    High Latency: Accessing large datasets stored in off-chip memory introduces delays that degrade system performance. Tensor operations, such as matrix multiplications and convolutions, require frequent memory accesses, and high latency in these operations can significantly reduce overall computation speed.
  2. 2.
    Limited Bandwidth: ML workloads involve the simultaneous transfer of large volumes of data, including weights, activations, and gradients. The limited bandwidth of traditional memory hierarchies often results in data transfer bottlenecks, slowing down the training and inference processes.
  3. 3.
    Inefficient Caching Strategies: Traditional caching mechanisms are not optimized for the structured and repetitive nature of tensor data. Without tensor-aware caching, caches are prone to evictions and cache misses, reducing data reuse and increasing latency.
  4. 4.
    Scalability Issues: As ML models and datasets grow in complexity, memory subsystems must scale to handle increasing data loads. Existing RISC-V architectures struggle to maintain performance when datasets exceed cache capacities, leading to frequent off-chip memory accesses and performance degradation.
To address these challenges, HERMES aims to design a high-performance memory hierarchy that integrates:
  • Shared L3 Caches: Enable efficient data sharing between CPU cores and accelerators, reducing latency and improving data availability.
  • Hybrid Memory Models: Combine DRAM and high-bandwidth memory (HBM) to balance capacity and speed, ensuring sufficient bandwidth for large-scale ML workloads.
  • Advanced Prefetching: Predict and pre-load data into caches to minimize latency and reduce the frequency of memory stalls.
  • Tensor-Aware Caching: Optimize cache replacement policies and data layouts specifically for tensor operations, increasing data reuse and reducing cache misses.
These solutions collectively aim to improve memory subsystem performance, enabling faster training and inference for complex ML models on RISC-V architectures.
IV Experimental Setup
Simulation Tools
To evaluate the HERMES memory hierarchy design, we utilize the following simulation tools:
  • Gem5: Gem5 is a versatile, cycle-accurate simulator that models CPU and memory interactions in computer architectures. It supports the RISC-V ISA and allows for detailed configuration of cache hierarchies, coherence protocols, and processor cores. We use Gem5 to simulate the entire memory subsystem and analyze performance metrics such as latency, bandwidth, and cache hit rates.
  • DRAMSim2: DRAMSim2 is a cycle-accurate simulator for DRAM memory systems. It provides detailed timing simulations for various DRAM technologies, including DDR3 and DDR4. DRAMSim2 is integrated with Gem5 to model off-chip memory accesses accurately and measure the impact of different memory hierarchies on system performance.
Architecture Design
The HERMES memory hierarchy integrates several advanced features to optimize performance for machine learning workloads. The key components of the architecture are as follows:
  • Shared L3 Cache: The shared L3 cache is designed to facilitate data sharing between CPU cores and the Gemmini accelerator. It employs fine-grained coherence protocols to ensure data consistency and minimize latency caused by cache invalidations. The L3 cache size and associativity are configurable to balance performance and area constraints.
  • Hybrid Memory Model: HERMES combines traditional DRAM with High-Bandwidth Memory (HBM). DRAM provides large storage capacity, while HBM delivers high data transfer rates. This combination allows for efficient handling of large datasets while ensuring that critical data can be accessed with low latency.
  • Tensor-Aware Caching: The cache hierarchy includes tensor-aware caching strategies that optimize data reuse patterns. Tensor data is stored in a way that reduces cache evictions and maximizes hit rates during tensor operations such as convolutions and matrix multiplications.
  • Advanced Prefetching: The prefetching mechanism anticipates data needs by analyzing access patterns during ML workloads. Both stride prefetching and machine learning-based prefetching are employed to minimize cache miss penalties and improve overall system throughput.
The following architecture diagram illustrates the core components of the HERMES memory hierarchy and their interconnections: [^1] This diagram shows the relationship between CPU cores, the shared L3 cache, the Gemmini accelerator, and the hybrid memory model (combining DRAM and HBM). It highlights the data flow and memory pathways designed to optimize performance for ML workloads.
Figure 1: HERMES Memory Hierarchy Architecture Diagram
Workloads
We evaluate the HERMES memory hierarchy using a suite of representative machine learning workloads, including:
  • Convolutional Neural Networks (CNNs): CNNs involve intensive matrix multiplications and convolution operations, which stress the memory bandwidth and caching mechanisms. Models such as ResNet and VGG are used for evaluation.
  • Recurrent Neural Networks (RNNs): RNNs require frequent access to previous states, placing a high demand on low-latency memory. We use LSTM (Long Short-Term Memory) and GRU (Gated Recurrent Unit) models to measure performance.
  • Transformer Models: Transformers involve large-scale attention mechanisms and matrix multiplications, resulting in high memory access patterns. BERT (Bidirectional Encoder Representations from Transformers) and GPT (Generative Pre-trained Transformer) models are used to test scalability and bandwidth utilization.
Performance Metrics
The following performance metrics are used to evaluate the HERMES memory hierarchy:
  • Latency (ns): The average memory access time, including cache hits and misses, is measured to evaluate how efficiently the memory subsystem handles data requests.
  • Bandwidth (GB/s): The rate at which data is transferred between memory and the processor/accelerator. Higher bandwidth indicates better handling of large data transfers.
  • Cache Hit Rate (%): The percentage of memory accesses that are served from the cache. Higher cache hit rates reduce latency and improve overall performance.
  • Energy Consumption (µJ/operation): The energy consumed per memory operation is measured to evaluate the efficiency of the memory subsystem.
Simulation Configuration
The experimental configuration used for the simulations is as follows:
  • Processor: 4-core RISC-V processor with in-order execution.
  • L1 Cache: 32 KB per core, 8-way set associative.
  • L2 Cache: 256 KB per core, 8-way set associative.
  • Shared L3 Cache: 8 MB, 16-way set associative.
  • Memory: Hybrid memory model combining 8 GB DRAM and 4 GB HBM.
  • Coherence Protocol: MESI (Modified, Exclusive, Shared, Invalid) protocol for cache coherence.
  • Simulation Tools: Gem5 for CPU and cache simulation, DRAMSim2 for DRAM modeling.
The configurations are designed to balance accuracy and simulation runtime while providing insights into the performance benefits of the HERMES memory hierarchy.
Results and Discussion
This section presents the performance evaluation of the HERMES memory hierarchy using simulations. The key metrics analyzed include latency, bandwidth, cache hit rates, and energy consumption. The results are compared against a baseline RISC-V system to highlight the improvements offered by HERMES.
Latency and Bandwidth Improvements
The latency and bandwidth of the HERMES architecture were evaluated and compared to a baseline RISC-V system. The results are summarized in Table I.
TABLE I: Latency and Bandwidth Comparison
Architecture Latency (ns) Bandwidth (GB/s)
Baseline RISC-V 120 25
HERMES (Shared L3 Cache) 95 35
HERMES (Prefetching) 85 40
HERMES (Tensor-Aware Caching) 80 42
Figure 2 shows a significant reduction in latency with the HERMES configurations, while Figure 3 highlights the bandwidth improvements due to the hybrid memory model.
Figure 2: Memory Latency Comparison: HERMES configurations reduce latency compared to the baseline RISC-V system. The shared L3 cache, advanced prefetching, and tensor-aware caching contribute to this improvement.
Figure 3: Bandwidth Utilization Comparison: The hybrid memory model in HERMES increases bandwidth significantly, supporting high data transfer needs of ML workloads.
Cache Hit Rate Improvements
Cache hit rates were evaluated to determine the effectiveness of tensor-aware caching in HERMES. Table II shows the results.
TABLE II: Cache Hit Rate Comparison
Architecture Cache Hit Rate (%)
Baseline RISC-V 60
HERMES (Shared L3 Cache) 75
HERMES (Prefetching) 80
HERMES (Tensor-Aware Caching) 90
Figure 4 illustrates the improvements in cache hit rates. Tensor-aware caching aligns data storage with tensor operation patterns, resulting in fewer cache misses.
Figure 4: Cache Hit Rate Comparison: Tensor-aware caching in HERMES improves cache hit rates by reducing data evictions and optimizing data reuse patterns.
Energy Consumption
Energy consumption is a critical factor for ML workloads. Table III shows the energy consumption per memory operation for different configurations.
TABLE III: Energy Consumption Comparison
Architecture Energy Consumption
(µJ/operation)
Baseline RISC-V 50
HERMES (Shared L3 Cache) 40
HERMES (Prefetching) 38
HERMES (Tensor-Aware Caching) 35
Figure 5 highlights the energy savings achieved by HERMES through efficient memory access strategies.
Figure 5: Energy Consumption Comparison: HERMES reduces energy consumption by minimizing off-chip memory accesses and employing efficient caching strategies.
Discussion
The simulation results clearly demonstrate the advantages of the HERMES memory hierarchy over the baseline RISC-V system:
  • Latency Reduction: HERMES reduces latency by up to 33% through shared L3 caches, advanced prefetching, and tensor-aware caching.
  • Bandwidth Improvement: The hybrid memory model increases bandwidth by up to 68%, supporting the high data transfer needs of ML workloads.
  • Cache Efficiency: Tensor-aware caching improves cache hit rates by 50%, reducing the need for expensive off-chip memory accesses.
  • Energy Efficiency: HERMES reduces energy consumption by up to 30%, making it suitable for energy-efficient ML computation.
These improvements highlight the effectiveness of integrating advanced memory design techniques in RISC-V architectures to meet the demands of modern ML workloads.
Acknowledgment
The author expresses sincere gratitude to Professor Ray Simar for his invaluable guidance, support, and encouragement throughout the development of this project. This work was completed as part of COMP 554: Computer Systems Architecture at Rice University. Professor Simar’s deep expertise in computer architecture and his commitment to student learning have been instrumental in shaping this project. His insights and feedback have significantly improved our understanding of memory systems and their optimization for machine learning workloads.